OpenStudy (anonymous):

Joe climbs a 5 foot ladder to the top of a slide. Find the length of the slide, to the nearest hundredth, if the distance from the base of the ladder to the bottom of the slide is 9 feet. Note: The ladder forms a right angle with the ground.

OpenStudy (anonymous):

The length of the slide is 22.50 feet.

OpenStudy (anonymous):

Right, we have 5 and 9 as the 2 legs of a right triangle. So we use Pythagorus to find the hypotenuse (third side) hyp ^2 = sqrt(5^2+9^2) = sqrt(106) = 10.2956 = 10.30 feet to nearest hundredth.

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

Anytime:-)